The company has a target debt-equity ratio of 0.8, a cost of equity of 14.0%, a cost of debt of 8.0%, and is subject to a 30% corporate tax rate. The company has never built and
operated a refining plant in the past, preferring to sell the mined materials to other companies for treatment and ultimate sale. Consequently, it regards this project as
requiring an increase in the firm’s cost of capital of +2%

a) Calculate the appropriate cost of capital that the company should apply to this project?
Answer #1

The debt-equity ratio = 0.8/1. Hence, the weightage of debt in the company's capital = 0.8/(0.8+1) = 44.44%. The weightage of equity is 100 - 44.44 = 55.56%

Hence, the old WACC (weighted average cost of capital) = [Wd *kd(1-marginal tax)] +[We*ke]

= [44.44%* 0.08(1-0.30)] +[55.56%*0.14] = 0.0249 + 0.1026 = 0.1275 or 12.75%

The new cost of capital for the new project = 12.75% + 2% = 14.75%